# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

### v5

@@ 5.4 Ablation Study
Figure 2: Effectiveness of GAT  explainers in identifying prediction-relevant graph structures using $\text{fidelity}^{+}$  (higher is better) and $\text{fidelity}^{-}$  (lower is better).
@@
To answer RQ2, Figure 2 compares the $\text{fidelity}^{+}$  and $\text{fidelity}^{-}$ , and Figure 3 compares the precision and recall of the explanations derived from SOTA methods (GNNExplainer, PGExplainer, and SubgraphX) with those derived from ProvExplainer. Later, we will analyze specific case studies in § 6.
$\text{fidelity}^{+}$  and $\text{fidelity}^{-}$ In the APT dataset from [mukherjee2023sec], ProvExplainer performs competitively for both $\text{fidelity}^{+}$  and $\text{fidelity}^{-}$ ProvExplainer and SubgraphX have competitive $\text{fidelity}^{+}$  performance and outperforming both GNNExplainer and PGExplainer. For $\text{fidelity}^{-}$  ProvExplainer performs the best by having the lowest score after k=10 but interestingly SubgraphX performed the worst which leads to the conclusion that SubgraphX is not identifying all the important subgraph structures, so when we remove graph structure that SubgraphX deemed unimportant it is still changing the prediction thats why in Figure 3 we can see it having the lowest recall value for GAT  network.
ProvExplainer was had the best performance across both fidelity types for DARPA TC  dataset. In the DARPA and APT datasets, the security-aware features of ProvExplainer provide a clear advantage in extracting important nodes from the provenance graphs whose removal changes the prediction, moreover low $\text{fidelity}^{-}$  indicate that the GNN model is particularly identifying those security-aware features since if we remove any other part of the graph the prediction does not change. For Fileless Malware  dataset ProvExplainer’s performance is moderate for $\text{fidelity}^{+}$ As we increased the size of $k$, ProvExplainer became the best explainer but for the lower values of $k$, SubgraphX performed well. But similar to performance of APT dataset from [mukherjee2023sec] we saw that SubgraphX only focusing on some particular structures and not accounting for all the important substructures that are important for prediction as indicative of a high $\text{fidelity}^{-}$
Across all the datasets, ProvExplainer shows an increase in fidelity values up to around $k=25$. ProvExplainer shows this improvement for $\text{fidelity}^{+}$  across the APT datasets. Unfortunately, this trend also appears in $\text{fidelity}^{-}$  but not much pronounced. This is because for long running and stealthy attacks, even the security-aware features are not enough, so it still missed some important graph structures that are important to the underlying GNN detection model. But, the increase of $\text{fidelity}^{-}$  of other explainers are more rapid when compared to ProvExplainer. Also, as $k$ values are increased we see an inverse trend for $\text{fidelity}^{+}$  and proportional trend for $\text{fidelity}^{-}$ , suggesting that when we either incorporate structures that produce unintended side effect and the label does not flip for $\text{fidelity}^{+}$  and for $\text{fidelity}^{-}$  we accidentally incorporate important subgraph structures whose removal changes the prediction.
Precision and Recall. In the APT dataset from [mukherjee2023sec], ProvExplainer demonstrates comitative precision performance for both GAT  and GraphSAGE  networks compared to other GNN explainers, although it achieves comparatively high recall. This suggests that ProvExplainer identifies more relevant instances, albeit at the cost of recall. In the DARPA TC dataset, ProvExplainer significantly outperforms other explainers in both precision and recall, respectively, for both GAT  and GraphSAGE  networks. This indicates that ProvExplainer is highly reliable in generating explanations for the APT datasets. In the Fileless Malware  dataset category, ProvExplainer achieves second best precision and recall, with values across GAT  and GraphSAGE , respectively. GNNExplainer performs the well in this category, since Fileless Malware  does not have intricate subgraph structures that are observed in APTs so GNNExplainer can identify them.
ProvExplainer’s performance highlights a pattern of excelling in the APT datasets, where it achieves both high precision and perfect recall, suggesting a strong alignment with the characteristics and structure of this dataset. Meanwhile, its performance on the Fileless Malware  dataset is moderate, and it has comparatively lower precision in the APT dataset from [mukherjee2023sec], where all explainers appear to face challenges. Regarding differences between GAT  and GraphSAGE  networks, ProvExplainer exhibits consistently higher performance in both precision and recall across these networks within each category. Other explainers, such as SubgraphX, show an improvement in precision when using GraphSAGE , although differences in recall are less pronounced. This pattern suggests that GraphSAGE  enhances the precision of some explainers without significantly impacting recall, although this effect is not uniform across all other SOTA explainers. GraphSAGE  focuses on the immediate neighbors of nodes, therefore it is less susceptible to irrelevant graph nodes.
A notable pattern emerges when comparing ProvExplainer against other explainers across dataset. ProvExplainer achieves high recall in the structured DARPA TC dataset, capturing relevant features more effectively than PGExplainer and GNNExplainer, which show much lower recall. Conversely, in the more complex APT dataset from [mukherjee2023sec], all explainers, including ProvExplainer, struggle to achieve high precision, suggesting that the dataset’s complexity or feature overlap may limit their discriminatory power. In the Fileless Malware  dataset, the generally high recall across explainers indicates their effectiveness in capturing indicative features, although precision remains challenging. These insights underscore ProvExplainer ’s strengths in structured datasets like DARPA TC and Fileless Malware  and its limitations in more complex APT scenarios like APT dataset from [mukherjee2023sec], highlighting the influence of TTPs used by attackers on explainer performance.
6 Case Studies
@@
To demonstrate ProvExplainer in a realistic setting, we analyze three case studies from the DARPA [darpa] datasets. In each study, we use an explanation size of 40 nodes and refer to the GAT  model. We list the most important features from the surrogate DT and qualitatively analyze the explanations from ProvExplainer and the SOTA GNN explainers. Detailed system-level analyses can be found in the appendix (§ A.5).
6.1 FiveDirections: Browser Extension
@@
Description. An attacker targets the Firefox browser using a malicious extension to deploy the drakon malware. The attacker writes drakon directly to disk and exploits a compromised browser extension masquerading as a password manager to execute malicious powershell code, gaining deeper access and control over the system, as illustrated in Figure 4.
Features: Degree centrality, pattern #1, and pattern #2.
System Interpretation. This attack graph contains the pattern #1 and pattern #2 in large numbers, which are often observed in the APT stages of deepen access and lateral movement. The pattern #1 is created when malwares have to transfer information among them selves. Therefore, after reading from a file, the malware updates the template and configuration files that other malwares in the system will read before executing their task. The pattern #2 is created since the malware in the system are copies therefore they need to read reads the same template files and .dll files as the original malware. Therefore, these patterns captures the dependency correlation among malware processes, where multiple instances of drakon exploit Firefox vulnerabilities via a malicious extension to spread to different parts of the system. This pattern is emblematic of malware processes cloning themselves to persist in the system as well as maintain operational integrity. Further, the pattern #2 highlights how malware processes move laterally to exfiltrate sensitive data and read configuration files. Since the attack graph contains high quantities of file read and write flowers for a relative small number of nodes, it leads to a high degree centrality.
ProvExplainer vs. SOTA explainers. 4(b) compares the explanations of ProvExplainer and those of SOTA GNN explainers, focusing on their efficacy in identifying security-aware elements. ProvExplainer excels by highlighting the malware replicating itself from template files and accessing sensitive system files. ProvExplainer isolates security-relevant structures in the graph, significantly increasing the end-user trust in the detection.
@@
6.2 FiveDirections: Copykatz
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #2.
(d) Pattern #5.
Description. In a sophisticated attack, a hijacked version of usdoj.gov exploits Firefox to deploy drakon malware to the victim host. Then, drakon uses the elevate driver to escalate privileges and masquerade as the runtimebroker system program. Finally, the malicious runtimebroker instance connects to a command and control (C2) server to download and execute Copykatz (an older version of Mimikatz) to harvest and exfiltrate host credentials, as illustrated in Figure 5.
Features: Degree centrality, pattern #2, and pattern #5.
System Interpretation. The attack graph highlights high degree centrality and two key patterns: #2 and #5. The pattern #2 captures the initial access by starting multiple malicious processes and writing to template files. The template files are written to by other malware child processes. This stage enables the Windows Application Programming Interface and cryptographic operations necessary for the malware’s functionality. Following this, the malware disguises itself as Firefox, and executes the Copykatz payload.
The malicious Firefox instance then distributes its payload through writing to temporary files, setting the stage for a subsequent malicious Firefox instance to trigger a flood of sensitive file reads and making network connections. This chain of actions, marked by inherited functionalities from Copykatz, forms the pattern #5. The replication of this malware leverages the same essential system libraries, indicating a meticulous design to maintain operational consistency throughout the malicious process chain. Similarly to the previous case study, due to high volume of network connections and processes creation the degree centrality is very high for the attack graph.
ProvExplainer vs. SOTA explainers. GNNExplainer and PGExplainer pinpoint the stage where multiple Firefox processes connect to the C2 servers. This activity traces back to the malicious runtimebroker instance. Notably, SubgraphX detects an alternate trajectory where a Firefox process, instead of reaching out to external servers, spawns another process aimed at local content manipulation, showcasing the malware’s versatility in engaging with both external and internal resources for its objectives.
While SOTA explainers have demonstrated proficiency in identifying the final stages of this data breach, only ProvExplainer effectively captured both the initial infection and its propagation. This distinction underscores the importance of recognizing early-stage indicators for root cause analysis.
6.3 Trace: Phishing E-mail
(a) Attack Graph.
(b) SOTA GNN explainations.
(c) Pattern #1.
(d) Pattern #2.
Description. The attacker first launches a phishing campaign to compromise the identity of an employee. Leveraging the employee’s identity, the attacker then targets other employees with deceptive emails containing links to a malicious website. This website installs a Trojan in the victims’ computers, which then creates multiple copies of itself, overflowing the system. These cloned Trojans read sensitive user files while the original Trojan achieves persistence in the system, as illustrated in Figure 6.
Features: Clustering coefficient, pattern #1 and pattern #2.
System Interpretation. In the Trace APT scenario, clustering coefficient, pattern #1 and pattern #2 elucidate the malware’s structure. Pattern #2 reveals the Trojan’s propagating activity, where it reads from template files and prepares to replicates itself to establish a foothold. By masquerading as benign programs, the malware disguises its malicious processes, allowing it to proliferate undetected. The clustering coefficient of the graph increases because of multiple malware reading the template files by forming large process creation and file read clusters.
The Trojan, after establishing itself, impersonates system processes to writing a malicious script, leading to other malicious process reading that script as seen in pattern #1. It illustrates the dependency correlation among the cloned malware instances, interacting with similar system configuration files to operate efficiently as well as constructing a detailed profile of the target system. Ultimately, a series of intra-malware communication happens where the malware reads from script and write to a file, and other malware update the script based on the current configuration. So, that the subsequent malware that read from the file gets the updated configuration.
ProvExplainer vs. SOTA explainers. GNNExplainer identifies the initial malware staging behaviors, including interactions with shared libraries and cache files. Meanwhile, PGExplainer captures the benign structure of xfce4-appfinder—a lightweight desktop environment for UNIX systems, invoked by a DARPA script to simulate an enterprise environment. SubgraphX identifies the malicious inheritance behavior of Firefox executing a template to generate multiple malware clones. However, it overlooks sensitive file accesses, a detail exclusively captured by ProvExplainer. ProvExplainer comprehensively traced the malware kill chain from payload deployment and clone creation to the final step of accessing sensitive files.
7 Discussion and Future Work
GNN Model Capacity. If decision trees trained with interpretable features provide accuracy comparable to GNN models, why do we need resource intensive GNN-based IDS? In an IDS context, the automatic adaptability and generality provided by GNN models are necessary for quickly understanding the execution profiles of a wide variety of programs. DTs have many properties that, while unproblematic in an explanation setting, are highly undesirable for security detectors. Beyond the difference in expressive power, DTs require labelled data, which can be provided by the GNN in an explanation setting, but will not be available for security detection. Further, DTs rely on features that are derived from known data, so they are less able to generalize to unseen data. Therefore, for accurate and robust security detection, GNN models are strongly preferred. The DT’s expressive power is weaker in comparison to GNNs, and demands human understanding as a prerequisite for adequate feature engineering.
Is ProvExplainer Model Agnostic? ProvExplainer is model agnostic and works on black-box provenance-based IDS. This line of research is not novel, as previous works  [jacobs2022ai], have used decision trees as surrogate models for deep neural networks before. ProvExplainer’s novelty stems from the development of interpretable graph features to enable surrogate DTs to provide explanations for decisions about heterogeneous graph data in the system provenance domain. The interpretation of advanced graph features in the system provenance domain enables domain experts to better understand the system-level evidence used by provenance-based IDS.
Capturing Subgraph Modifications. To address the weakness of whole-graph aggregation washing out the influence of small malicious subgraph modifications, ProvExplainer’s DT explanations can be combined with general-purpose graph explanations techniques which were originally not suitable for the security domain, but can be tied to problem space actions by ProvExplainer’s interpretable features. In future work, we will combine ProvExplainer’s security-aware results that contain global perspective with the exploration stage of SubgraphX to better assist in identifying important subgraph structures.
@@ 7 Discussion and Future Work
Adversarial Manipulation of Graph Features. What are the implications of attackers using knowledge of the graph structural features to design adversarial attacks against the explainer? The explainer models used in ProvExplainer is separate from the black-box detection model. Fooling both the interpretable surrogates and the underlying black-box model is the ideal case for an attacker seeking to avoid suspicion, and an attacker who can attack the explainer this way can also attack the detection model directly [goyal2023sometimes, mukherjee2023sec]. Creating robust detection and explanation systems that can withstand adversarial manipulation is a critical open research problem, and is beyond the scope of this work.
Reliable Explanation for Unreliable GNN Models. Recent studies have made significant strides in investigating evasive attack generation against provenance-based ML detectors [goyal2023ndss, provninja2023sec], a noteworthy perspective that raises questions about the robustness and stability of the ML models in question. However, the core underlying assumption driving model explainer research — that the stability and trustworthiness of ML models can always be assured — may not necessarily hold. In future work, we plan to investigate the efficacy of model explainers under adversarial conditions in ML models could enhance their reliability, providing insights into their performance in real-world adversarial scenarios and contribute to the development of robust systems.
Behavioral Query Language for System Provenance. In the future work we hope to develop a graph query language using defined behavioral features to support efficient retrieval of provenance records based on program behaviors. The graph query will be based on the graph features, allowing the operator to search for programs using high level descriptions of their behavior, which will be converted into graph features in the query engine. The direction will also benefit significantly from advancements in subgraph mining algorithms. Future works may consider specialized provenance storage systems to optimize data retrieval.
Support for Fine-Grained Detection Tasks. Recent developments in provenance-based security detection systems have trended towards fine-grained node and edge level anomaly detection [rehman2024flash, zengy2022shadewatcher]. The explanation requirements of these detectors differ from those of whole-graph anomaly detectors and classifiers. In contrast to using explanations to narrow down the context for consideration by security practitioners, fine-grained models will need explainers to bring in relevant context to aid human analysis. Providing security-aware explanation support to advanced fine-grained security detectors is an exciting and important direction for future work.
8 Conclusion

### v7

@@ VI-D Graph Structural Feature Evaluation
Figure 4: Comparison of GAT explainers in identifying important structures using $\text{fidelity}^{+}$ (higher is better) and $\text{fidelity}^{-}$ (lower is better).
@@
VI-E Ablation Study
Figure 3 shows the contributions of different features to the overall surrogate DTs agreement of approximating a GAT model. Interested readers may find this study with GraphSAGE in the appendix, Figure 8. We separate the features into two groups: subgraphs and graph structural features. We evaluate each group individually and then we look at the their combined effect. We further separate the group into individual features (e.g., top-$K$ subgraphs) and look at the effect of each constituting features in the group. Furthermore, to explore the performance of different groups, we evaluate different graph structural features individually and collectively.
In the APT dataset, the top-$K$ subgraphs individually show different levels of agreement for both the Enterprise and Supply-Chain subsets, indicating that distinct subgraphs capture complementary signals. A similar trend is observed for the DARPA TC and Fileless Malware datasets, although the increase in agreement levels plateaus as $K$ approaches 4. This suggests that the highest-ranked subgraphs encode the most informative patterns, supporting our claim that the decision path extracted from the surrogate DT effectively captures the discriminative power of the GNN.
We examine individual structural features, certain metrics (i.e., eigenvector centrality and clustering coefficient) consistently exhibit high agreement in the APT datasets, particularly in DARPA TC (Trace and Theia) dataset. In contrast, other metrics (i.e., closeness and betweenness centrality) show low agreement across most APT datasets. This behavior arises because, in the context of APT attacks, centrality measures often encode information that misleads the model: benign graphs contain values similar to those of attack graphs. As attackers commonly leverage system processes [darpa] to do their bidding and system processes exhibit similar structural characteristics, which reduces their discriminative power.
In the APT dataset, where attackers employ stealthy techniques, which suggests the necessity of integrating subgraph structures with graph structural features to construct a comprehensive feature set as combined features provide stronger alignment than relying on individual features. ProvExplainer achieves the best performance across all datasets when all features are combined.
An interesting pattern is that the Fileless Malware dataset consistently performs exceptionally well, particularly with individual features like eigenvector centrality and clustering triangles, which achieve near-perfect alignment scores. This suggests that the structural properties in the Fileless Malware dataset are distinctively different compared to the benign programs they are masquerading.
VI-F ProvExplainer vs. SOTA Explainers
@@ VI-E Ablation Study
To answer RQ2, Figure 4 compares the $\text{fidelity}^{+}$ $\text{fidelity}^{-}$ and Figure 5 compares the precision/recall of the explanations derived from ProvExplainer and SOTA methods (GNNExplainer [gnnexplainer], PGExplainer [pgexplainer], and SubgraphX [subgraphx]). We analyze specific case studies in § VII.
$\text{Fidelity}^{+}$ and $\text{Fidelity}^{-}$. In the APT dataset from [46], ProvExplainer achieves the best performance for both $\text{fidelity}^{+}$ and $\text{fidelity}^{-}$ as seen in Figure 4. ProvExplainer and SubgraphX show competitive $\text{fidelity}^{+}$ results, outperforming both GNNExplainer and PGExplainer. For $\text{fidelity}^{-}$ , GNNExplainer performs the worst, as it fails to identify all the critical subgraph structures. Therefore, when graph structures deemed unimportant are removed, the predictions still change, resulting in high $\text{fidelity}^{-}$ value. ProvExplainer’s superior performance in the DARPA and APT datasets is attributed to its security-aware features, which provide a significant advantage in identifying and extracting critical nodes from provenance graphs. These nodes, when removed, lead to prediction changes, demonstrating ProvExplainer’s ability to capture essential graph components.
Additionally, ProvExplainer achieves high coverage of important nodes, reflected in its low $\text{fidelity}^{-}$ values, removing unimportant nodes does not affect predictions. In the Fileless Malware dataset, ProvExplainer demonstrates moderate performance at lower $k$ values (e.g., the total number of explanation nodes considered). However, as $k$ increases, ProvExplainer becomes the best performer in $\text{fidelity}^{+}$ , with SubgraphX ranking second. These results highlight that SubgraphX focuses on a limited number of structures and fails to provide comprehensive coverage of all the critical substructures essential for accurate predictions.
ProvExplainer’s $\text{fidelity}^{-}$ performance is hindered in the Fileless Malware dataset due to the unique substructures created by Fileless Malware attacks, which are not adequately captured by its security-aware features. If distinguishing subgraphs present in the test set are absent from the training set, these critical structures are missed, leading to gaps in the explanations. However, the $\text{fidelity}^{-}$ performance of other explainers declines more rapidly as $k$ increases in the Fileless Malware dataset. This suggests that general-purpose GNN explainers, which operate by calculating masks around edges and subgraphs, perform better when unique subgraph structures are present for each specific attack. It is important to note that ProvExplainer’s primary objective is not to serve as a zero-day vulnerability detector but rather to produce human-interpretable explanations.
Precision and Recall. In the APT datasets, ProvExplainer demonstrates superior precision and recall performance for both GAT and GraphSAGE networks compared to other GNN explainers as seen in Figure 5. This highlights ProvExplainer’s reliability in generating accurate explanations for the APT datasets by identifying relevant subgraph structures that include nodes mentioned in the ground truth. A similar trend is observed in Fileless Malware when explaining GAT , although ProvExplainer achieves the second-best precision and recall in this case. These results emphasize ProvExplainer’s strength in explaining complex APT datasets, where common TTPs from the MITRE framework are frequently reused (i.e., similar subgraphs are formed). However, ProvExplainer struggles with the Fileless Malware dataset, where unique structures in the testing set are absent in the training set, illustrating the impact of novel TTPs on explainer performance.
GNNExplainer performs well in the Fileless Malware dataset category due to its fidelity-driven approach, where GNNExplainer calculates masks for substructures for every instance, enabling it to effectively identify diverse structures. Similarly, SubgraphX shows improved precision and recall when explaining GraphSAGE compared to GAT , as SubgraphX focuses on identifying important subgraphs based on fixed-size node neighborhoods. These neighborhoods influence the message-passing algorithm of GraphSAGE , ensuring that the explanations generated by SubgraphX align closely with GraphSAGE ’s computational behavior.
ProvExplainer’s performance demonstrates its strength in the APT and Fileless Malware datasets, achieving both high precision and perfect recall. This suggests a strong alignment with the inherent characteristics and structure of these datasets. However, in the more complex APT dataset [46], all explainers, including ProvExplainer, face challenges in maintaining high precision. This limitation arises due to the dataset’s complexity and the novel TTPs employed by APT attackers during certain stages. These TTPs are represented in only a subset of the graphs. Consequently, when these specific graphs appear exclusively in the testing set, ProvExplainer lacks the corresponding subgraphs for explanation, resulting in missed patterns and a subsequent decline in precision.
Table II: Explanation actionability comparison between ProvExplainer and SOTA GNN explainers measured by human actionability metric. Absolute human actionability metric (top) and relative factor versus ProvExplainer (italic, bottom). Lower values reflect better actionable.
Dataset
GNNExplainer PGExplainer SubgraphX ProvExplainer
DARPA APT
Dataset [darpa]
3.29
[…]
2.60
[…]
2.36
[…]
2.18
[…]
APT Dataset
from [46]
1.97
[…]
1.71
[…]
1.85
[…]
1.24
[…]
Fileless Malware
from [47]
2.04
[…]
1.74
[…]
1.77
[…]
1.34
[…]
VI-G Human Actionability Study
[…]
VII Case Studies
@@
To demonstrate ProvExplainer in a realistic setting, we analyze case studies from the DARPA [darpa] datasets. In each study, we use an explanation size of 40 nodes and refer to the GAT model. We list the most important features from the surrogate DT and qualitatively analyze the explanations from ProvExplainer and the SOTA GNN explainers. One more case study (§ A-E) and detailed analyses in the appendix (§ A-F).
VII-A FiveDirections: Browser Extension [darpa]
@@
Description. An attacker targets the Firefox browser using a malicious extension to deploy the drakon malware. The attacker writes drakon directly to disk and exploits a compromised browser extension masquerading as a password manager to execute malicious powershell code, gaining deeper access and control over the system (Figure 6).
Features: Degree centrality and pattern #1.
System Interpretation. The pattern shown in 6(c) provides a detailed representation of the initial stages (i.e., initial compromise and privilege escalation) of APT attack lifecycle orchestrated by the Drakon APT, showcasing a sequence of file and process interactions. Process P1 (pass_mgr.exe) serves as the initial pivot point, leveraging File F1 (firefox.exe) to gain access to the system through a compromised browser backdoor. This initial stage involves P1 writing to multiple files, including File F2 (ntdll.dll), File F3 (sysmain.sdb), File F4 (passwordfile.dat), and File F5 (RICHED20.DLL). These operations signify the deployment of malicious shellcode, preparation for privilege escalation, and establishment of a foothold within the system.
The transition to Stage 2 is marked by the activities of Process P2 (pass_mgr.exe), which reads from File F5 (RICHED20.DLL) and writes to critical files such as File F3 and File F4. These interactions indicate the progression to privilege escalation and credential harvesting, leveraging tools like Copykatz and Drakon APT’s elevate driver. The repeated use of the same process executable (pass_mgr.exe) across different stages points to sophisticated APT TTPs such as process injection or executable reuse.
Since the attack graph contains high quantities of file read and write flowers for a relatively small number of nodes between these flowers, it leads to a high degree of centrality, which means a high amount of process interaction between static artifacts like files. The graph’s structural measures complement the subgraph’s interpretation and identifies the TTPs employed by Drakon APT.
ProvExplainer vs. SOTA explainers. 6(b) compares the explanations of ProvExplainer and those of SOTA GNN explainers, focusing on their efficacy in identifying security-aware elements. ProvExplainer excels by highlighting the malware replicating itself from template files and accessing sensitive system files. ProvExplainer isolates security-relevant structures in the graph, significantly increasing the end-user trust in the detection.
@@
VIII Discussion and Future Work
Limitations. ProvExplainer depends on previously observed graph patterns, as the surrogate DT is trained on discriminant subgraph patterns mined from an existing dataset. Therefore, ProvExplainer is unlikely to perform well on attacks that contain previously unseen subgraph patterns. We recognize the inherent limitations of our approach and, therefore, set explicit boundaries for ProvExplainer’s interpretive range, aiming to validate detection results and provide actionable intelligence for known attacks. Preconditions that may prevent ProvExplainer from explaining some attacks: (1) the detection model must detect the attack, (2) the attack graph contains distinct subgraphs identified during the mining phase, and (3) the attack’s distinctive subgraphs are small enough to be feasibly pattern-matched.
@@ VIII Discussion and Future Work
Adversarial Manipulation of Graph Features. The explainer models used in ProvExplainer is separate from the black-box detection model. Fooling both the surrogates and the detection model is the ideal case for an attacker seeking to avoid suspicion, and an attacker who can attack the explainer this way can also attack the detection model directly [goyal2023sometimes, 46]. Creating robust explanation framework that can withstand adversarial manipulation is a critical open research problem, and is beyond the scope of this work.
IX Conclusion